# VLD: Visual Language Goal Distance for Reinforcement Learning Navigation

### v1

@@ Appendix I Ordinal Consistency: Qualitative Results
Besides geometric overlap noise (Section 3.2.2), we experiment with one more distance-noising method. Since recurrent policies (LSTMs, GRUs) can inherently filter out i.i.d. Gaussian noise [lstm, gru], we do not use this most straightforward approach for applying noise. Instead, we introduce temporally correlated noise based on Ornstein–Uhlenbeck (OU) process, a type of “red noise” widely used in RL exploration [ddpg, noise_test]. For each episode, we maintain a noise state updated as:
@@
Most failures arise from trajectories where the agent becomes stuck exploring visually similar but spatially incorrect regions. Equipping the agent with explicit memory—either short- or long-term—of previously visited locations, for example via a topological map or episodic graph [vint, feng2025imagegoalnavigationusingrefined], would likely mitigate such failures (see Figure 19(a)). However, memory augmentation is outside the scope of this work, as our focus is on evaluating VLD as a standalone navigation signal.
@@
A third failure mode originates from the properties of the distance function itself. In these near-match failures, the agent encounters a view that appears deceptively similar to the goal image even though the true goal is still physically distant (Figures 19(c)19(d)). However, because the policy relies solely on a scalar distance signal—which is naturally biased toward feature overlap—the agent may incorrectly conclude that it has reached the goal and stop prematurely.

### v2

@@ Appendix I Ordinal Consistency: Qualitative Results
Besides geometric overlap noise (Section 3.2.2), we experiment with one more distance-noising method. Since recurrent policies (LSTMs, GRUs) can inherently filter out i.i.d. Gaussian noise [15, 5], we do not use this most straightforward approach for applying noise. Instead, we introduce temporally correlated noise based on Ornstein–Uhlenbeck (OU) process, a type of “red noise” widely used in RL exploration [24, 16]. For each episode, we maintain a noise state updated as:
@@
Most failures arise from trajectories where the agent becomes stuck exploring visually similar but spatially incorrect regions. Equipping the agent with explicit memory—either short- or long-term—of previously visited locations, for example via a topological map or episodic graph [41, 8], would likely mitigate such failures (see Figure 20(a)). However, memory augmentation is outside the scope of this work, as our focus is on evaluating VLD as a standalone navigation signal.
@@
A third failure mode originates from the properties of the distance function itself. In these near-match failures, the agent encounters a view that appears deceptively similar to the goal image even though the true goal is still physically distant (Figures 20(c)20(d)). However, because the policy relies solely on a scalar distance signal—which is naturally biased toward feature overlap—the agent may incorrectly conclude that it has reached the goal and stop prematurely.
@@
Figure 21: Goal images used in the real-world navigation experiments.
Appendix M Real-World Experiments
Table 12: Real-world navigation results. Success rate (SR) and average number of steps required to reach each goal. Each goal is evaluated with three repetitions. The human operator trajectory is reported as a reference baseline.
Method G1 G2 G3 G4 G5 Total
Human (steps) 29 14 36 28 13
VLD + (GeoNoise) SR (↑) 1.00 1.00 1.00 0.67 1.00 0.93
VLD + (GeoNoise) Avg. Steps (↓) 58.6 138.3 249.3 342.5 80.7
FGPrompt-EF SR (↑) 0.00 0.33 0.67 1.00 1.00 0.60
FGPrompt-EF Avg. Steps (↓) 326.0 271.5 75.3 106.3
M.1 Experiment Setup
We evaluate the proposed navigation approach on a physical TurtleBot4-Light robot [32] equipped with a Logitech MX Brio 705 RGB camera providing a $90^{\circ}$ field of view. The camera is mounted at a height of approximately $1\,\text{m}$ to match the RL training setup. Compass and GPS observations required by the policies are estimated from the TurtleBot4 odometry.
We define five goal locations, each represented by a goal image captured at the target position (Figure 21). The initial position ($S$) for the first goal ($G_{1}$) is fixed such that the goal location is not visible at the start of the episode, requiring the agent to explore before reaching the target. For subsequent goals, the starting position corresponds to the previous goal location (i.e., $G_{2}$ starts at $G_{1}$, $G_{3}$ starts at $G_{2}$, and so on). Each goal is evaluated with three independent repetitions, resulting in $15$ total navigation attempts per method.
The navigation tasks vary in difficulty. Some goals require only limited exploration, in which the agent can observe elements of the goal scene after a small rotation or short movement (e.g., $G_{1}\rightarrow G_{2}$). Others require exploring the environment in the correct direction before the goal becomes visible (e.g., $S\rightarrow G_{1}$, $G_{4}\rightarrow G_{5}$), while the most challenging tasks involve navigating across multiple spaces or rooms before detecting the goal location (e.g., $G_{2}\rightarrow G_{3}$, $G_{3}\rightarrow G_{4}$).
For reference, we also record trajectories executed by a human operator controlling the robot. The operator is familiar with the environment and knows the location of each goal beforehand; therefore, these trajectories do not involve exploration and represent an approximate lower bound on the number of steps required to reach each target. Each forward action corresponds to approximately $0.5\,\text{m}$ of motion. The resulting number of steps required to reach each goal is reported in Table 12. A video of the human-driven trajectories is provided in the supplementary material.
M.2 Results
We report success rate (SR) as the number of successful goal reaches divided by the total number of attempts per agent ($15$ attempts in total). The average number of steps is computed over successful runs only.
Table 12 shows that our method successfully reaches every goal at least twice across the three trials, achieving an overall success rate of $0.87$, while FGPrompt-EF achieves $0.60$. We observe distinct navigation behaviors between the two approaches. Similar to its behavior in simulation, our agent periodically performs $360^{\circ}$ rotations to evaluate how the predicted visual-language distance changes across directions before committing to forward motion. While this behavior increases SPL in simulation due to additional rotations, it proves beneficial in the real-world setting where perceptual cues may be subtle or appear from viewpoints different from the goal image. By repeatedly reassessing the environment from multiple viewpoints, the agent can detect weak goal cues and gradually move toward regions where the predicted distance decreases.
In contrast, FGPrompt-EF tends to select a navigation direction early and pursue it with limited reassessment of alternative directions. When the chosen direction is incorrect, the agent often struggles to recover and redirect toward the correct path. This behavior is particularly evident for goal $G_{1}$, where the policy frequently terminates with a stop action far from the target in locations where the current view contains no visible cues from the goal image. We hypothesize that this behavior may partially result from sim-to-real discrepancies affecting perception or stopping decisions.
Interestingly, FGPrompt-EF performs well on $G_{4}$, where the agent consistently selects the correct direction early in the episode. In such cases, committing to a single direction allows the agent to reach the goal efficiently. Nevertheless, across most tasks the exploration-driven strategy of our method proves advantageous, even requiring fewer steps on average for three of the five goals.
To further illustrate the behavior of our method, Figure 22 presents a representative navigation trajectory in the real-world environment. We also provide additional videos of trajectories in the supplementary material.
Finally, we emphasize that these real-world experiments represent a small-scale evaluation constrained by available hardware and experimental time. While the results demonstrate promising transfer of the learned visual-language distance representation from simulation to the physical robot, larger-scale evaluations across diverse environments and longer navigation tasks will be required to fully characterize the robustness and generalization capabilities of the approach.
(a) After initial exploration with no visible goal cues, the agent chooses to explore the hallway.
(b) While the predicted distance initially increases, the agent detects a slight decrease when turning toward the left, indicating a potentially promising direction.
(c) Moving further along the hallway reveals visual cues consistent with the goal image, causing the predicted distance to drop sharply.
(d) The agent follows the decreasing distance signal and successfully navigates to the goal.
Figure 22: Example real-world navigation trajectory using VLD. Even when the goal object is not directly visible, the predicted visual-language distance provides directional guidance that allows the agent to progressively move toward the target.